OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 13 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of leaves in the two spirals?

OpenStudy (anonymous):

fibonnaci sequence 1,1,2,3,5,8,13,21, 34,55,...

OpenStudy (anonymous):

therefore since one is 13 more than other, one is 21 and one is 34

OpenStudy (anonymous):

do u want me to show u the choices cause thats not what i have its fraction stuff

OpenStudy (anonymous):

@mak_12: Could you do a little bit calculation here: x / 13 = 1.6 => x = 13 * 1.6 = 20.8 ( ~21)